OpenStudy (anonymous):

A circle is described by the equation x2 + y2 + 14x + 2y + 14 = 0. What are the coordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

(-7, -1), 36 units (7, 1), 36 units (7, 1), 6 units (-7, -1), 6 units

OpenStudy (anonymous):

the question with the answers

OpenStudy (anonymous):

((7,1) with 6units..

OpenStudy (anonymous):

it was wrong

OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)?

OpenStudy (anonymous):

x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

OpenStudy (anonymous):

just complete there squares...and then see what it comes to be...

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation?

OpenStudy (anonymous):

(x − 1)2 − (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 3 (x + 1)2 + (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 1

OpenStudy (anonymous):

its standard form.. (x-a)2 + (y-b)2 = r2

OpenStudy (anonymous):

also visit here for clearence... http://www.mathsisfun.com/algebra/circle-equations.html